# Learning Piecewise Residuals of Control Barrier Functions for Safety of Switching Systems using Multi-Output Gaussian Processes

### v1

@@ I Introduction
Many real-world control problems are modeled as hybrid dynamical systems, which involve a coupling between continuous dynamics and discrete events. Research in the stability analysis of such systems has grown significantly in recent decades [5, 9, 16]. More recently, Gaussian processes (GPs) are used to approximate hybrid residual dynamics model [10]. However, in the modern context of control applications, the focus has expanded from mere stability assurance to encompass safety verification. CBFs have recently provided a powerful theoretical tool for synthesizing controllers that ensure the safety of dynamical systems. They were initially applied to autonomous driving and bipedal walking [3]. Stability guarantees are then added via a unified control Lyapunov function-control barrier function (CLF-CBF) framework by solving a state-dependent quadratic programming (QP) problem at each time step to compute the control input [20]. In the context of hybrid systems, barrier functions were initially used as safety certificates in [19]. After the introduction of CBFs, CLF-CBF design has been extensively used in safety-critical methods for hybrid systems [3, 17, 18]. However, there are very few studies on learning-based safety verification of switching systems via CBFs in the presence of uncertainty.
@@
Learning-based CBFs for uncertain systems can be divided into two approaches. The first approach learns the unknown dynamics and uses the approximated model to derive safety certificates. In [23], a GP is employed to learn the unmodeled dynamics for the safe navigation of quadrotors with CBFs. A GP-based systematic controller synthesis method is introduced in [14], offering an inherently safe control strategy. In [11], the authors employ a Bayesian approach to learn model uncertainty while ensuring safety during the learning process. However, learning the entire system dynamics using GP is computationally expensive. As a result, a prevalent assumption is to learn each element of the vector fields individually and overlook their correlations. Also, all the aforementioned works do not consider uncertainty in the actuation term to reduce the complexity of the learning-based model.
@@
These limitations lead us to the second approach, which approximates the projection of the residual onto the CLF and CBF. In [22], an episodic learning method using neural networks is employed to approximate the impact of unmodeled dynamics on system safety, eliminating the need for exhaustive data collection. This idea has been implemented on the hybrid bi-pedal robot in [8]. However, methods employing neural networks must achieve precise approximations of the dynamics to formally ensure safety. In general, this requirement is challenging to meet, but GPs can provide uncertainty quantification with analytical confidence bounds. In [6], a GP-based min-norm controller stabilizing an unknown control affine system using CLFs is introduced. The effect of the uncertainty in the stability constraint is approximated by a GP model. The same method is applied to CBF design in [7]. This technique results in the approximation of a real-valued function instead of a vector field, leading to a significant reduction in the model complexity for high-dimensional systems. In [1], the effect of uncertainty on high-order CBFs is quantified using GPs.
@@ II-C Impact of uncertainty in the switching systems
In this paper, our primary focus is characterizing the adverse effects of uncertainty on the proposed safety-critical control method. We then adopt a supervised learning approach to make our design robust to piecewise residuals of the form (8). Thus, a dataset that contains information about piecewise residuals in all regions is needed.
@@ III-B Batch multi-output GP for piecewise residuals
Since $k_{r}^{i}(\boldsymbol{x},\boldsymbol{x}^{\prime})$’s are positive definite kernels, then by the definition, there is a feature map $\varphi(\boldsymbol{x})$ such that $k_{r}^{i}(\boldsymbol{x},\boldsymbol{x}^{\prime})=\varphi^{T}_{i}(\boldsymbol{x})\varphi_{i}(\boldsymbol{x}^{\prime})$ for $i\in\{1,\dots,m+1\}$. Now, Let $y^{i}$ and $y^{\prime i}$ be the $i^{th}$ element of the corresponding vectors. We have
@@ III-D Feasibility analysis
We can easily verify that $\beta_{r}^{2}\Sigma_{r}^{h}=\begin{bmatrix}\beta_{r}\boldsymbol{l}_{r}^{1}&A_{r}^{h}(\boldsymbol{x})\end{bmatrix}^{T}\begin{bmatrix}\beta_{r}\boldsymbol{l}_{r}^{1}&A_{r}^{h}(\boldsymbol{x})\end{bmatrix}$ and $M_{r}/1=S_{r}(\boldsymbol{x})$. We know that $\Sigma_{r}^{h}(\boldsymbol{x})$ is symmetric positive-definite (SPD). So, $M_{r}$ is also symmetric and we can characterize its definiteness by Schur complement theorem. Since its lower right block is positive definite and its corresponding Schur complement $M_{r}/\beta_{r}^{2}\Sigma_{r}^{h}(\boldsymbol{x})>0$ by our assumption, we can conclude that $M_{r}$ is also positive definite.
@@ III-E Data collection and model fitting
We use an episodic data collection method. In the first episode, we run an initial roll-out using the CLF and CBF designed based on the nominal system. Using nominal system (7), we run QP controller (6) with the nominal values of the time derivative of $V(\boldsymbol{x})$ and $h(\boldsymbol{x})$. Then, we record the system’s trajectory until it reaches an unsafe state. Next, we partition the resulting dataset into $R$ datasets $\mathcal{D}_{r}$ as described in Section III-B. Subsequently, we fit MOGPs configured with the kernel functions $k_{r}$ to batches $\mathcal{D}_{r}$, for $r=1,\dots,R$. To use (18) and (19) for prediction, we need to infer the hyperparameters $\boldsymbol{\theta}_{r}$ for each individual kernel $k_{r}^{i},i\in\{1,\dots,m+1\}$ in the structure of $k_{r}$. Each set of parameters is obtained by minimizing the negative log marginal likelihood as
@@ IV Simulation results
The distance between the ego car and the front car is denoted by $z$. The front car is moving at a constant velocity $v_{0}=10\hskip 2.0ptm/s$. Thus, the distance between the cars is governed by $\dot{z}=v_{0}-x_{2}$, where we have the initial distance $z(t_{0})=z_{0}$. The control objective is to reach a target speed $v_{d}=18\hskip 2.0ptm/s$ while maintaining a safe distance to the vehicle in front.
We consider the Lyapunov function $V(\boldsymbol{x})=(x_{2}-v_{d})^{2}$ to stabilize the velocity at the desired value. Also, we consider CBF $h(\boldsymbol{x})=z-T_{h}x_{2}$, where $T_{h}=1.6$ is a predefined threshold to maintain a safe distance from the car in front proportional to the current velocity.
We selected infinitely differentiable squared exponential kernel [24], for $k_{i},i=1,2$ base kernels in the configuration of $k_{r}$. We used the process in Section III-E to collect the dataset and optimize the batch MOGP model hyperparameters, which collected a total number of $390$ data samples in the initial roll-out and a total of $999$ data samples in all episodes. The simulation is started from the initial state $\boldsymbol{x}_{0}=\begin{bmatrix}0&14\end{bmatrix}^{T}$ and $z_{0}=140\hskip 2.0ptm$. The ego car starts in the region $\mathcal{R}_{1}$ from a safe distance behind the front car. First, the controller increases the velocity to converge to the desired speed resulting in a decrease in $z$. Then, the ego car transitions into region $\mathcal{R}_{2}$ and approaches the front car which triggers the activation of the safety constraint (23c). Consequently, the controller must reduce the velocity to keep the safe distance. Due to the adverse effect of the piecewise residuals, the Single GP-SOCP and nominal QP controller cannot maintain the safe distance. However, the proposed MOGP-SOCP controller effectively avoids this unsafe behavior by leveraging the batch MOGP design. A simulation video showing the performance of the nominal and MOGP-SOCP controller is provided. Figure 1 shows a snapshot of a simulation when all controllers result in the closest distance to the front car. It can be verified that only MOGP-SOCP controllers can keep the safe distance during the simulation.
@@
In Figure 2, we aim to compare the performance of the proposed method with the baseline and nominal controllers. For your reference, we provided the trajectories of the Oracle true design. It has been illustrated in Figure 2 that the single GP-SOCP controller did not converge to the desired velocity, while MOGP-SOCP and nominal QP controllers could reach $v_{d}$ in the first $4$ seconds. Then, the ego car transitioned to $\mathcal{R}_{2}$ at $t=6\hskip 2.0pts$ which affects the rate of change of the velocity. As it approaches the front car, the controllers must decrease the velocity in response to remain safe. The CBF $h(\boldsymbol{x})$ and control input $u$ are also illustrated in this figure. This plot shows that the single GP-SOCP and nominal QP controllers violate the safety condition approximately at $t=15\hskip 2.0pts$ and $t=17\hskip 2.0pts$, respectively. However, the MOGP-SOCP controller could maintain the safe distance to the front car during the simulation. Also, it can be verified from the control input trajectory that the MOGP-SOCP controller could recover the true system’s performance. However, the single GP-SOCP controller does not generate a smooth control input which is a result of poor uncertainty quantification between the switching times.

### v2

@@ I Introduction
Many real-world control problems are modeled as hybrid dynamical systems, which involve a coupling between continuous dynamics and discrete events. Research in the stability analysis of such systems has grown significantly in recent decades [7, 10, 18]. More recently, Gaussian processes (GPs) are used to approximate hybrid residual dynamics model which can be integrated into a stabilizing model predictive control design ensuring probabilistic constraint satisfaction [11]. However, in the modern control applications, the focus has expanded from mere stability assurance to encompass safety verification. CBFs have recently provided a powerful theoretical tool for synthesizing controllers that ensure the safety of dynamical systems. They were initially applied to autonomous driving and bipedal walking [3]. Stability guarantees are then added via a unified control Lyapunov function-control barrier function (CLF-CBF) framework by solving a state-dependent quadratic programming (QP) problem at each time step to compute the control input [22]. In the context of hybrid systems, barrier functions were initially used as safety certificates in [21]. After the introduction of CBFs, CLF-CBF design has been extensively used in safety-critical methods for hybrid systems [3, 19, 20]. However, there are very few studies on learning-based safety verification of switching systems via CBFs in the presence of uncertainty.
@@
Learning-based CBFs for uncertain systems can be divided into two approaches. The first approach learns the unknown dynamics and uses the approximated model to derive safety certificates. In [24], a GP is employed to learn the unmodeled dynamics for the safe navigation of quadrotors with CBFs. A GP-based systematic controller synthesis method is introduced in [15], offering an inherently safe control strategy. In [12], the authors employ a Bayesian approach to learn model uncertainty while ensuring safety during the learning process. Learning the complete system dynamics with GPs is computationally demanding. Therefore, a common approach is to learn each component of the vector fields individually, disregarding their correlations. Furthermore, existing studies typically neglect uncertainty in the actuation term to simplify the learning-based model.
@@
These limitations lead us to the second approach, which approximates the projection of the residual onto the CLF and CBF. In [9], an episodic learning method using neural networks is employed to approximate the impact of unmodeled dynamics of hybrid bi-pedal robot on system safety, eliminating the need for exhaustive data collection. However, methods employing neural networks must achieve precise approximations of the dynamics to formally ensure safety. In general, this requirement is challenging to meet, but GPs can provide uncertainty quantification with analytical confidence bounds. In [8], a GP-based min-norm controller stabilizing an unknown control affine system while guaranteeing safety using CLF-CBF method is introduced. The effect of the uncertainty in the stability and safety constraints is approximated by a GP model. This technique results in the approximation of a real-valued function instead of a vector field, leading to a significant reduction in the model complexity for high-dimensional systems. In [1], the effect of uncertainty on high-order CBFs is quantified using GPs.
@@
A conference version of this paper will appear in [2]. This extended version includes additional details omitted in the conference paper.
@@ II-C Impact of uncertainty in the switching systems
In this paper, our primary focus is characterizing the adverse effects of uncertainty in the proposed safety-critical control method. We then adopt a supervised learning approach to make our design robust to piecewise residuals of the form (8). Thus, a dataset that contains information about piecewise residuals in all regions is needed.
@@ III-B Batch multi-output GP for piecewise residuals
Since $k_{r}^{i}(\boldsymbol{x},\boldsymbol{x}^{\prime})$’s are positive definite kernels, then by the definition, there is a feature map $\varphi(\boldsymbol{x})$ such that $k_{r}^{i}(\boldsymbol{x},\boldsymbol{x}^{\prime})=\varphi^{T}_{i}(\boldsymbol{x})\varphi_{i}(\boldsymbol{x}^{\prime})$ for $i\in\{1,\dots,m+1\}$ [4]. Now, Let $y^{i}$ and $y^{\prime i}$ be the $i^{th}$ element of the corresponding vectors. We have
@@ III-D Feasibility analysis
We can easily verify that $\beta_{r}^{2}\Sigma_{r}^{h}=\begin{bmatrix}\beta_{r}\boldsymbol{l}_{r}^{1}&A_{r}^{h}(\boldsymbol{x})\end{bmatrix}^{T}\begin{bmatrix}\beta_{r}\boldsymbol{l}_{r}^{1}&A_{r}^{h}(\boldsymbol{x})\end{bmatrix}$ and $M_{r}/1=S_{r}(\boldsymbol{x})$. We know that $\Sigma_{r}^{h}(\boldsymbol{x})$ is symmetric positive-definite (SPD). So, $M_{r}$ is also symmetric and we can characterize its definiteness by Schur complement theorem. Since its lower right block is positive definite and its corresponding Schur complement $M_{r}/\beta_{r}^{2}\Sigma_{r}^{h}(\boldsymbol{x})>0$ by our assumption, based on [6], we can conclude that $M_{r}$ is also positive definite.
@@ III-E Data collection and model fitting
We use an episodic data collection method. In the first episode, we run an initial roll-out using the CLF and CBF designed based on the nominal system. Using nominal system (7), we run QP controller (6) with the nominal values of the time derivative of $V(\boldsymbol{x})$ and $h(\boldsymbol{x})$. Then, we record the system’s trajectory until it reaches an unsafe state. Next, we partition the resulting dataset into $R$ datasets $\mathcal{D}_{r}$ as described in Section III-B. Subsequently, we fit MOGPs configured with the kernel functions $k_{r}$ to batches $\mathcal{D}_{r}$, for $r=1,\dots,R$. To use (18) and (19) for prediction, we need to infer the hyperparameters $\boldsymbol{\theta}_{r}$ for each individual kernel $k_{r}^{i},i\in\{1,\dots,m+1\}$ in the structure of $k_{r}$. Each set of parameters is obtained by minimizing the negative log marginal likelihood [17] as
@@ IV Simulation results
The distance between the ego car and the front car, denoted by $z$, is governed by the equation $\dot{z}=v_{0}-x_{2}$, where $v_{0}=10\hskip 2.0ptm/s$ represents the constant velocity of the front car. The control objective is to achieve a target speed of $v_{d}$ while ensuring a safe distance from the front vehicle. To stabilize the velocity at the desired value, a Lyapunov function $V(\boldsymbol{x})=(x_{2}-v_{d})^{2}$ is considered. Additionally, a CBF $h(\boldsymbol{x})=z-T_{h}x_{2}$, with $T_{h}=1.6$, is utilized to maintain a safe distance proportional to the current velocity.
We selected infinitely differentiable squared exponential kernel [25], for $k_{i},i=1,2$ base kernels in the configuration of $k_{r}$. We used the process in Section III-E to collect the dataset and optimize the batch MOGP model hyperparameters, which collected a total number of $390$ data samples in the initial roll-out and a total of $999$ data samples in all episodes. The simulation is started from the initial state $\boldsymbol{x}_{0}=\begin{bmatrix}0&14\end{bmatrix}^{T}$ and $z_{0}=140\hskip 2.0ptm$. The ego car starts in the region $\mathcal{R}_{1}$ from a safe distance behind the front car. First, the controller increases the velocity to converge to the desired speed resulting in a decrease in $z$. Then, the ego car transits into region $\mathcal{R}_{2}$ and approaches the front car which activates the safety constraint (23c). Consequently, the controller must reduce the velocity to keep the safe distance. Due to the adverse effect of the piecewise residuals, the Single GP-SOCP and nominal QP controller cannot maintain the safe distance. However, the proposed MOGP-SOCP controller effectively avoids this unsafe behavior by leveraging the batch MOGP design. A simulation video showing the performance of the nominal and MOGP-SOCP controller is provided. Figure 1 shows a snapshot of a simulation when all controllers result in the closest distance to the front car. It can be verified that only MOGP-SOCP controllers can keep the safe distance during the simulation.
@@
In Figure 2, we aim to compare the performance of the proposed method with the baseline and nominal controllers. For your reference, we provided the trajectories of the oracle true design. It has been illustrated in Figure 2 that the single GP-SOCP controller did not converge to the desired velocity, while MOGP-SOCP and nominal QP controllers could reach $v_{d}$ in the first $4$ seconds. Then, the ego car transited to $\mathcal{R}_{2}$ at $t=6\hskip 2.0pts$ which affects the rate of change of the velocity. As it approaches the front car, the controllers must decrease the velocity in response to remain safe. The CBF $h(\boldsymbol{x})$ and control input $u$ are also illustrated in this figure. This plot shows that the single GP-SOCP and nominal QP controllers violate the safety condition approximately at $t=15\hskip 2.0pts$ and $t=17\hskip 2.0pts$, respectively. However, the MOGP-SOCP controller could maintain the safe distance to the front car during the simulation. Also, it can be verified from the control input trajectory that the MOGP-SOCP controller could recover the true system’s performance. However, the single GP-SOCP controller does not generate a smooth control input which is a result of poor uncertainty quantification between the switching times.